OpenStudy (kweb2017):

Given that (12,3) is on the graph of f(x), find the corresponding point for the function 8f(x).

OpenStudy (kweb2017):

(12,24)

OpenStudy (badsheep):

Let g(x)= 8f(x), find the point (x, g(x)) the only x coordinate I'm given is 12, hence it,s (12,g(12))=(12,8*f(12))=(12,(8*3))=(12,24)